OpenStudy (anonymous):

Estimate using compatible numbers. seventy-eight percent of one hundred ninety-six

OpenStudy (anonymous):

It is about 160. You see, you can use the method: \[\frac{ is }{ of } = \frac{ percent }{ 100 }\] The of is 196 since it is next to the word 'of'. And the percent is 78%. Now plug it in.\[\frac{ is }{ 196 } = \frac{ 78 }{ 100 }\] You can solve for is by cross multiply. So 196 x 78 is equal to 15288. Divide it by 100 to get 152.88, which is close to about 160.